# Data-driven decision-making under uncertainty with entropic risk measure

### v1

@@ 6 Distributionally robust insurance pricing
Since the insurer does not know the true distribution ${\mathbb{P}}$, and the marginal distribution ${\mathbb{P}}_{i}^{\prime}s$ are also unknown to the homeowners, decisions based on the SAA can be optimistically biased. To address this, the insurer solves the following distributionally robust insurance pricing problem, which minimizes the worst-case entropic risk:
@@
$\mathop{\rm min}_{(z_{i},\pi_{i})_{i=1}^{M}}\mathop{\rm sup}_{\mathbb{Q}\in\mathcal{B}_{\infty}(\epsilon)}\quad$ $\rho_{\mathbb{Q}}\left(\sum_{i=1}^{N}z_{i}\xi_{i}-\sum_{i=1}^{N}\pi_{i}\right)$ (19a)
subject to $\rho_{\hat{{\mathbb{P}}}_{i,N}}\left(\xi_{i}-z_{i}\xi_{i}+\pi_{i}\right)\leq\rho_{\hat{{\mathbb{P}}}_{i,N}}\left(\xi_{i}\right)$ $\forall i\in[M],$ (19b)
@@
where $\mathbb{Q}$ lies in the type-$\infty$ Wasserstein ambiguity set given in (12). From Theorem 2, the above problem can be reformulated as the following regularized exponential cone program:
@@
$\mathop{\rm min}_{(z_{i},\pi_{i})_{i=1}^{M}}\quad$ $\rho_{\hat{{\mathbb{P}}}_{N}}\left(\sum_{i=1}^{N}z_{i}\xi_{i}-\sum_{i=1}^{N}\pi_{i}\right)+\epsilon\|z\|_{*}$ (20)
subject to $\rho_{\hat{{\mathbb{P}}}_{i,N}}\left(\xi_{i}-z_{i}\xi_{i}+\pi_{i}\right)\leq\rho_{\hat{{\mathbb{P}}}_{i,N}}\left(\xi_{i}\right)$ $\forall i\in[M].$ (21)
@@ 6.1 Numerical Experiments
To illustrate our model and algorithm using numerical experiments, we consider $5$ homeowners with risk aversion parameters $\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\alpha_{5}=2.9,2.7,2.5,2.3,2.1$. For all methods, except SAA and Opt, we use K-fold cross-validation procedure given in Algorithm 4 with $K=5$ to construct the validation dataset $\mathcal{X}$ of loss scenarios. Let $\hat{\rho}_{\epsilon}$ denote the empirical entropic risk of loss scenarios in $\mathcal{X}$ for a given radius $\epsilon$. The selection of $\epsilon$ varies across methods as follows:
  • Traditional CV approach: The radius is chosen based solely on $\hat{\rho}_{\epsilon}$.
  • BS-EVT: A two-component GMM is fitted to the loss scenarios in set $\mathcal{X}$ and the radius is selected based on bias-corrected values of $\hat{\rho}_{\epsilon}$
  • BS-Match: Entropic risk matching is employed to fit a GMM to the loss scenarios in $\mathcal{X}$ and the radius is selected based on bias-corrected values of $\hat{\rho}_{\epsilon}$.
  • Opt: The radius is calibrated based on the entropic risk over the test data-points.
After choosing the radius $\epsilon$ for each model, we find the optimal decision by solving SAA problem using all the training samples. The distribution of losses is modeled using a Gaussian copula with Gamma-distributed marginals and a correlation coefficient $r$ across $M$ homeowners. Let $\Sigma\in\mathbb{S}_{+}^{M\times M}$ denote the positive-definite covariance matrix:
$\Sigma=r\mathbf{1}\mathbf{1}^{\top}+(1-r)I,$
where $\mathbf{1}$ is a vector of all ones and $I$ is the identity matrix. Samples are generated by first drawing a multivariate normal vector $\omega\sim\mathcal{N}(\mathbf{0},\Sigma)$, and transforming $\omega$ element-wise using the standard normal cdf, $\varsigma=\Phi_{0,1}(\omega)$. Finally, the losses $\xi$ are sampled from a Gamma distribution:
$\xi_{i}=\Gamma^{-1}(\varsigma,\kappa_{i},\lambda_{i}),$
where $\Gamma^{-1}$ is the inverse cdf of the Gamma distribution, and $\kappa_{i}$ and $\lambda_{i}$ are its shape and scale parameters, respectively.
The boxplots are obtained for $500$ repeated experiments, each experiment done by sampling $N$ training samples from the Gaussian copula. The regularizer $\epsilon$ is chosen from a set containing 00 and $20$ equally-spaced values in the interval $[10^{-7},1]$ on the log scale. Using the values of correlation coefficient and parameters of marginal distribution, we sample $100,000$ i.i.d. data points from the Gaussian copula that are used to evaluate the out-of-sample performance.
Figure 5: Variation in the insurer’s out-of-sample entropic risk as a function of the number of training samples $N$. This analysis considers the scenario where each homeowner observes samples from different $\Gamma$ marginals.
Figure 6: Variation in the insurer’s out-of-sample entropic risk as a function of the number of training samples $N$. This analysis considers the scenario where each homeowner observes samples from the same $\Gamma$ marginals.
Homeowners have different marginal distribution.
In this experiment, the scale and location parameters of the $\Gamma$-distributed marginals of 5 homeowners are given by $(8,0.4)$, $(8.5,0.425)$, $(9,0.45)$, $(9.5,0.475)$, and $(10,0.5)$, respectively. The correlation coefficient is set to $0.5$.
Figure 5 shows the variation in the out-of-sample entropic risk with the number of training data-points. It is evident from the figure that our models, BS-EVT and BS-Match,outperform the traditional CV method and the SAA, resulting in a lower out-of-sample entropic risk. This suggests that both the EVT and entropic risk matching based approaches offer better generalization and robustness compared to standard empirical risk minimization techniques when applied to insurance risk estimation under limited training data scenarios.
@@
Figure 7: Variation in the premium per unit expected coverage of the insurer with number of training samples $N$.
Figure 8: Variation in the insurer’s out-of-sample entropic risk as a function of the correlation coeffficient $r$.
Homeowners have same marginal loss distribution
To analyze the impact of the training sample size, $N$, and the correlation coefficient, $r$, on the out-of-sample risk and premium per unit expected coverage, we assume the marginal distribution of loss across households follows $\Gamma(10,0.5)$.
@@
Figure 9: Variation in the premium per unit expected coverage of the insurer with correlation coefficient $r$.
As depicted in Figure 6, out-of-sample risk decreases with increase in the number of training data points. This decrease occurs because the premium per unit coverage better reflects the true risk exposure of the homeowners as shown in Figure 7. Consequently, with larger values of $N$, the insurer can extract higher premium from the homeowners for the same coverage level. The average out-of-sample loss, denoted as $\bar{\xi}$ in Figure 7 remains constant across households.
Figure 8 demonstrates that the out-of-sample entropic risk initially increases with the correlation coefficient $r$, eventually stabilizing. A similar trend is observed in Figure 9 which shows the variation in the average premium per unit expected coverage across homeowners. This behavior can be attributed to the insurer setting higher premiums, potentially leading homeowners to opt out of insurance coverage.
@@ 7 Conclusions
In this paper, we propose two practical approaches to mitigate the bias inherent in the empirical entropic risk estimator, which arises when the empirical distribution is used in place of the true distribution in entropic risk calculations. We prove that the bias-corrected entropic risk estimator converges almost surely to the true risk. However, when a GMM is fitted using maximum likelihood estimation and the bias is estimated via bootstrapping, the bias-corrected estimator often underestimates the true risk in finite samples. Our proposed approaches address this underestimation problem by estimating the bias in samples drawn from a “bias-aware” GMM fitted to the data.
To address the optimistic bias associated with optimizing the entropic risk measure, we propose to use type-$\infty$ Wasserstein ambiguity sets in DRO models. This is because other type-$p$ Wasserstein ambiguity sets with $p<\infty$ lead to unbounded worst-case entropic risk for any decision. For piecewise-concave loss functions, we provide equivalent reformulations of DRO problems as finite-dimensional convex optimization problems using Fenchel duality. Additionally, we introduce a distributionally robust version of a well-known insurance pricing problem. Numerical experiments demonstrate that our bias correction procedures help identify the appropriate radius for the ambiguity set, resulting in premium pricing and coverage policies that achieve lower out-of-sample entropic risks when compared to cross-validation methods that do not account for bias in the validation scenarios.
Appendix A Proofs
@@ A.1 Proof of Lemma 1
We also have that:
@@
$\ \geq\ -2G/C$
Hence that ${\mathbb{E}}[\exp(\alpha\ell({{\boldsymbol{\xi}}}))]\geq\exp(-2G/C)$
@@
Hence,
$0\leq\text{Var}[\exp(\alpha\ell({{\boldsymbol{\xi}}}))]={\mathbb{E}}[\exp(2\alpha\ell({{\boldsymbol{\xi}}}))]-{\mathbb{E}}[\exp(\alpha\ell({{\boldsymbol{\xi}}}))]^{2}\leq{\mathbb{E}}[\exp(2\alpha\ell({{\boldsymbol{\xi}}}))]\leq\frac{2G}{C-2}.$

### v3

@@ 6 Distributionally robust insurance policy
As discussed in the previous section, decisions based on the empirical data can be optimistically biased. To address this, the insurer solves the following distributionally robust insurance pricing problem, which minimizes the worst-case entropic risk:
@@
$\begin{array}[]{lll}\mathop{\rm min}&\displaystyle\mathop{\rm sup}_{\mathbb{Q}\in\mathcal{B}_{\infty}(\epsilon)}\;\rho^{\alpha_{0}}_{\mathbb{Q}}\left(\boldsymbol{z}^{\top}{{\boldsymbol{\xi}}}-{{\boldsymbol{1}}}^{\top}\boldsymbol{\pi}\right)\\ \text{s.t.}&\boldsymbol{\pi}\in\mathbb{R}_{+}^{M},\,\boldsymbol{z}\in[0,1]^{M}\\ &\rho^{\alpha_{h}}_{\hat{{\mathbb{P}}}_{h,N}}\left(\pi_{h}+(1-z_{h})\xi_{h}\right)\leq\rho^{\alpha_{h}}_{\hat{{\mathbb{P}}}_{h,N}}\left(\xi_{h}\right)&\forall h\in[M],\end{array}$
@@
where $\mathbb{Q}$ lies in the type-$\infty$ Wasserstein ambiguity set given in (11). Since the loss function is linear, it follows from Corollary 6 that the problem can be reformulated as the following regularized exponential cone program:
@@
$\begin{array}[]{lll}\mathop{\rm min}&\rho^{\alpha_{0}}_{\hat{{\mathbb{P}}}_{N}}\left(\boldsymbol{z}^{\top}{{\boldsymbol{\xi}}}-{{\boldsymbol{1}}}^{\top}\boldsymbol{\pi}\right)+\epsilon\|{{\boldsymbol{z}}}\|_{*}\\ \text{s.t.}&\boldsymbol{\pi}\in\mathbb{R}_{+}^{M},\,\boldsymbol{z}\in[0,1]^{M}\\ &\rho^{\alpha_{h}}_{\hat{{\mathbb{P}}}_{h,N}}\left(\pi_{h}+(1-z_{h})\xi_{h}\right)\leq\rho^{\alpha_{h}}_{\hat{{\mathbb{P}}}_{h,N}}\left(\xi_{h}\right)&\forall h\in[M].\end{array}$ (19)
The proposed setting is motivated by Bernard et al., (2020) who assumes that both the insurer and households are expected utility maximizers, with complete information on the true loss distributions for each household and their risk aversion parameters. Our approach relaxes the assumption of known loss distributions by providing only samples of the loss distribution to both the insurer and the households. While Bernard et al., (2020) impose some additional assumptions to analytically characterize pricing and coverage decisions under partially correlated risks, our data-driven method formulates the problem as a tractable exponential cone program (18) which needs to be solved numerically. Additionally, the insurer’s robustified (DRO) problem (19) retains this tractability. At optimality, the constraints hold with equality due to the monotonicity of entropic risk measure, that is, $\rho^{\alpha_{h}}_{\hat{{\mathbb{P}}}_{h,N}}\left(\pi_{h}+(1-z_{h})\xi_{h}\right)=\rho^{\alpha_{h}}_{\hat{{\mathbb{P}}}_{h,N}}\left(\xi_{h}\right)$ which can be equivalently written as:
$\pi_{h}=\frac{1}{\alpha_{h}}\log\left(\frac{{\mathbb{E}}_{\hat{{\mathbb{P}}}_{h,N}}[\exp(\alpha_{h}\xi_{h})]}{{\mathbb{E}}_{\hat{{\mathbb{P}}}_{h,N}}[\exp(\alpha_{h}(1-z_{h})\xi_{h})]}\right).$ (20)
In Example 1, we computed the optimal premium price under full coverage $z_{h}=1$ in which case the premium in (20) equals the entropic risk. Furthermore, the demand response model (17), which links premiums to coverage, accounts for the asymmetry in risk perception between households—who rely on ^$\hat{\mathbb{P}}_{h,N}$—and the insurer, who uses ^$\hat{\mathbb{P}}_{N}$. This flexible framework adapts to households’ evolving responses as additional information is incorporated through observed catastrophic events, thereby updating the empirical distributions. It is worth noting that our study could easily be adapted to accommodate alternative demand response models as long as the insurers valuations of the insurance is a concave function of coverage. Moreover, the regulatory constraints enforcing minimum coverage requirements for each household could easily be integrated.
@@ 6.1 Numerical Experiments
The numerical experiments conducted in this section demonstrate the effectiveness of our proposed distributionally robust insurance pricing model under various conditions. Our main goal is to evaluate how different calibration methods for the radius $\epsilon$ influence the insurer’s out-of-sample entropic risk and investigate the structure of the optimal policies $(\boldsymbol{z},\boldsymbol{\pi})$ offered to the households. Each household’s loss follows a Gamma distribution $\Gamma(\kappa_{h},\lambda_{h})$, with shape $\kappa_{h}$ and scale $\lambda_{h}$ parameters specific to each household $h\in[M]$. The correlation among the losses of different households is modelled using a Gaussian copula, with $\Sigma=r\mathbf{1}\mathbf{1}^{\top}+(1-r)I$, where $r$ controls the amount of correlation among the different households, $\mathbf{1}$ is a vector of all ones, and $I$ is the identity matrix.
In the following experiments, we consider $M=5$ households with risk aversion parameters $\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\alpha_{5}=2.9,2.7,2.5,2.3,2.1$ and set the insurer’s risk aversion parameter to $\alpha_{0}=2$. In Algorithm 5, we select the radius of the ambiguity set $\epsilon$ from set $\mathcal{E}$ which contains twenty equally-spaced values in the interval $[0,6]$. To generate an instance, we sample $N$ loss scenarios for each household and evaluate the out-of-sample performance, by generating $10^{6}$ i.i.d. data points (test data). This procedure is repeated over $100$ instances to obtain the statistics presented in the subsequent sections.
We consider five different calibration methods. Models BS-Match and BS-EVT use the $5$-fold CV Algorithm 5 together with Algorithm 2 and Algorithm 3 within the bootstraping procedure, respectively, i.e., $K=5$. CV model corresponds to the traditional $5$-fold CV where we set $\delta=0$ in step 4 of Algorithm 5. The model labeled as Oracle uses the test data for calibrating the radius $\epsilon$ at the validation step. Finally, model SAA solves problem (18) and does not involve any calibration.
All experiments described in the following sections were conducted in Python. The MOSEK 10.1 solver was used to solve exponential cone programs, while the entropic risk matching was performed on a GPU using the POT library (Flamary et al.,, 2021).
6.1.1 Case with mild correlation
(a) Insurer’s out-of-sample entropic risk.
(b) Optimal radius $\epsilon^{*}$ selected by different methods.
Figure 5: Comparison of the effects of training sample size $N$ on out-of-sample entropic risk (left) and optimal radius $\epsilon^{*}$ (right). Boxplots present the statistics after $100$ resampling of datasets, and diamonds presents the mean for each $N$.
In the first experiment, we examine the effect of sample size on the out-of-sample risk observed by the insurer. In the base scenario, all households have a common Gamma-distributed marginal loss distribution, $\Gamma(10,0.45)$, with a correlation coefficient $r=0.5$. This configuration allows us to focus on the effects of sample size $N$ and correlation coefficient $r$ on the insurer’s decisions, controlling for variability from differing marginal distributions. Similar insights hold for cases with heterogeneous marginal distributions among households, discussed in Appendix B.6. The results are summarized in Figures 5-8. Figure 5(a) shows that both BS-Match and BS-EVT consistently outperform CV and SAA across different sample sizes, while achieving an out-of-sample entropic risk similar to the oracle-based calibration. This can be explained by looking at the ambiguity radius $\epsilon^{*}$ chosen by each method. In Figure 5(b), we observe that CV typically chooses $\epsilon^{*}$ values that are significantly lower than the optimal choice, while the BS-Match and BS-EVT choices are closer to optimal. This discrepancy results from CV’s estimation procedure, which underestimates the true entropic risk for each $\epsilon$ (as discussed in Section 4), leading to select an overly optimistic $\epsilon^{*}$ in step 7 of Algorithm 5. This can be seen in Figure 6 where we plot the variation in the entropic risk estimator with the radius $\epsilon$ for each model with $N=1000$ (similar plots are obtained in Appendix B.7 for $N\in\{500,5000,10000\}$). In contrast, BS-Match and BS-EVT better estimate the trend in the variation of true entropic risk with $\epsilon$, thus making a more informed choice for $\epsilon^{*}$.
Figure 6: Statistics of entropic risk estimators for different radius and $N=1000$ after $100$ resampling of datasets
(a) Insurer’s estimate of the optimal entropic risk.
(b) Average optimal premium per unit of expected coverage.
@@ 6.1.1 Case with mild correlation
Figure 7: Comparison of the effects of number of training samples $N$ on insurer’s estimate of optimal entropic risk and average premium per unit of expected coverage.
Figure 7(a) depicts the estimation of the optimal entropic risk computed by each method. The results have a similar interpretation as Figure 4 in Section 4.3. We observe a significant underestimation of the true entropic by CV and SAA, while the estimation of the BS-Match and BS-EVT stay close to the estimation of the optimal calibration Oracle. We observe that as the sample size increases, the optimal risk decreases and so is the case of each method’s optimal risk estimation. This is because it is optimal for households to pay higher premiums due to the increase in their estimate of the risk of their respective loss.
Figure 8: Effect of the number of training samples on the proportion of instances where premium exceeds twice the expected coverage for BS-Match. Households become less risk averse as we go from the left of the panel to the right one.
Consequently, the insurer can charge higher premiums per unit expected coverage as a function of $N$, see Figure 7(b) where the average premium per unit expected coverage across the $5$ households is given $(1/5)\sum_{h=1}^{5}\pi_{h}/(\bar{\xi}_{h}z_{h})$ with $\bar{\xi}_{h}$ as the expected loss of household $h$. We observe an increase in this ratio as $N$ increases, which can be explained from the observation both the insurer and households become more capable of accurately estimating their true risk, thus enabling the insurer to extract higher premiums from households for the same coverage level. Moreover, CV and SAA charge lower premium per unit of expected coverage compared to our proposed approaches because they underestimate the risk. Namely, both methods are overly optimistic regarding how to correct the estimation error due to sampling, effectively using an $\epsilon$ that is too low. To illustrate the variation in the households’ willingness to pay for insurance as a function of number of training samples, we use as a proxy, the proportion of instances where the premium is at least twice the expected coverage. For BS-Match, Figure 8 illustrates that highly risk-averse households pay higher premiums per unit of expected coverage more frequently even when the number of training samples is low. Additionally, as $N$ increases, the proportion approaches $100\%$ for all households.
(a) Insurer’s out-of-sample entropic risk.
(b) Optimal radius $\epsilon^{*}$ selected by different methods.
Figure 9: Comparison of the effects of correlation coefficient $r$ on out-of-sample entropic risk (left) and optimal radius $\epsilon^{*}$ (right).
(a) Insurer’s estimate of the optimal entropic risk.
(b) Average optimal premium per unit of expected coverage.
@@
Figure 10: Comparison of the effects of correlation coefficient $r$ on estimate of optimal entropic risk and premium per unit expected coverage.
(a) BS-Match.
(b) SAA.
Figure 11: Effect of correlation on the coverage proportion offered to households averaged across $100$ instances. Each panel represents a different household. Households become less risk-averse as we go from the left panel to the right one.
6.1.2 Effect of correlations
In the second experiment, we fix the sample size at $N=1000$. Similar to the previous experiment, we assume that all households share a common marginal loss distribution modeled by a Gamma distribution, $\Gamma(10,0.45)$. However, in this experiment, we vary the parameter $r$, which controls the pairwise correlation of losses between households. The correlation coefficient ranges from 00 (independent losses) to $1$ (comonotone losses). Figures 9-12 summarize the results. First, note that similar insights about the effectiveness of our proposed approaches can also be observed in this setting. Indeed, Figure 9(b) shows that both CV and SAA are over-optimistic, leading them to select lower $\epsilon^{*}$ values than those chosen by our proposed approaches. With regards to the behavior in terms of $r$, Figure 9(a) demonstrates that the out-of-sample entropic risk initially increases with the correlation coefficient $r$, but eventually stabilizes. This trend is intuitive: higher correlation among households’ losses means that extreme loss events are more likely to occur simultaneously, increasing the insurer’s risk exposure. Figure 10(a) shows the estimates of the optimal entropic risk produced by each model. BS-EVT and BS-Match overestimate the optimal entropic risk of the insurer compared to Oracle at low correlation levels. Figure 10(b) shows that the average optimal premium per unit of expected coverage also increases with the correlation coefficient. This reflects the insurer’s response to higher risk by charging higher premiums to compensate for the increased likelihood of large, simultaneous payouts. However, there is a diminishing return effect; beyond a certain point, further increases in correlation do not lead to significantly higher premiums per unit of expected coverage. As the correlation between household losses increases, the benefits of risk pooling diminish, so the insurer reduces coverage levels significantly to reduce the risk exposure, see Figure 11(a) and 11(b). While Figures 11(a) and 11(b) show that households receive more coverage with SAA than BS-Match, this high coverage exposes the insurer to higher risks than the optimal coverage in the case of highly correlated losses. Figure 12 demonstrates that as the correlation $r$ across households increases, the proportion of instances with premiums exceeding twice the expected coverage also increase, and this effect is more pronounced for more risk-averse households. The high risk-averse households secure greater coverage by paying high premiums to reduce their risk exposure.
Figure 12: Effect of correlation $r$ on proportion of instances where premium exceeds two times the expected coverage. Each panel represents a different household. Households become less risk averse as we go from the left panel to the right one.
@@ 7 Conclusions
In this paper, we propose two practical approaches to mitigate the bias inherent in the empirical entropic risk estimator, which arises when the empirical distribution is used in place of the true distribution in entropic risk calculations. We propose to correct the bias using bootstrapping and prove that such a bias-corrected entropic risk estimator converges almost surely to the true risk. However, when bootstrapping is based on a GMM fitted using MLE, the bias-corrected estimator continues to underestimate the true risk with finite samples. Our proposed approaches mitigate this underestimation problem by estimating the bias in samples drawn from a “bias-aware” GMM fitted to the data.
To address the associated optimistic bias when optimizing the entropic risk measure, we propose to use type-$\infty$ Wasserstein ambiguity sets in DRO models. For piecewise-concave loss functions, we provide equivalent reformulations of DRO problems as finite-dimensional convex optimization problems using Fenchel duality. We show that traditional K-fold CV used to tune the radius of the ambiguity set can result in the underestimation of risk and the proposed approaches can mitigate the underestimation problem. Specifically, we introduce a distributionally robust insurance pricing problem with a risk-averse insurer and risk-averse homeowners whose demand for insurance depends on their own empirical distribution of losses. Numerical experiments demonstrate that our bias mitigation procedures help identify the appropriate radius for the ambiguity set, resulting in premium pricing and coverage policies that achieve lower out-of-sample entropic risks when compared to CV methods that do not account for bias in the validation scenarios. The numerical results also validate the effectiveness of our DRO framework in improving insurance pricing strategies under uncertainty. The experiments underscore the critical role of accurate risk estimation in insurance pricing, especially for catastrophic events exacerbated by climate change. Insurers must consider not only the empirical data but also the uncertainty inherent in limited samples and the potential for extreme losses. Our findings suggest that incorporating distributional robustness into the pricing model allows insurers to balance the trade-off between premiums charged and coverage offered for potential losses. This approach can help mitigate the financial vulnerability exemplified by programs like the NFIP, which faces substantial deficits due to under-priced premiums and low uptake.
There are several avenues for extending this research. First, the proposed bias correction procedures can be adapted to estimate other risk measures, such as Conditional Value at Risk. Importantly, our methods do not depend heavily on the specific form of the entropic risk measure, aside from the fact that tail scenarios exert disproportionate influence on it. Exploring the effectiveness of these bias corrections in multi-stage settings, where entropic risk measures are widely used, such as control theory and reinforcement learning, would be valuable. Furthermore, investigating the potential benefits of bias correction in other contexts–such as portfolio optimization and inventory management–represents a promising direction for future research. Finally, leveraging advancements in solving exponential cone programs (see Ye and Xie,, 2021; Chen et al., 2024a, ) to address large-scale entropic risk minimization problems, potentially with integer constraints, offers a valuable opportunity for enhancing computational efficiency.
Acknowledgments
The first author’s research is supported by the NSERC, Canada, grant RGPIN-2024-05067. This research is supported by compute resources provided by Calcul Quebec and the Digital Research Alliance of Canada. We thank Tianyu Wang for his helpful comments on an earlier version of the paper.
Appendix A Additional proofs and results
@@ A.1 Proof of Lemma 1
[…]
where first equality follows from representing expectation of a random variable using its cumulative distribution distribution (cdf), the last inequality follows from Assumption 1 and $C>2$ is used to obtain the final result.
From Jensen’s inequality, we also have that:
@@
$\ =\ -2G/C,$
where the last inequality follows from Assumption 1. Hence, we have that ${\mathbb{E}}[\exp(\alpha\ell({{\boldsymbol{\xi}}}))]\geq\exp(-2G/C)$.
@@
where we used $C>2$ to obtain the last equality, see Assumption 1. Hence,
$0\leq\text{Var}(\exp(\alpha\ell({{\boldsymbol{\xi}}})))={\mathbb{E}}[\exp(2\alpha\ell({{\boldsymbol{\xi}}}))]-{\mathbb{E}}[\exp(\alpha\ell({{\boldsymbol{\xi}}}))]^{2}\leq{\mathbb{E}}[\exp(2\alpha\ell({{\boldsymbol{\xi}}}))]\leq\frac{2G}{C-2}.$